OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

jimthompson5910 (jim_thompson5910):

\[\Large -2(x-3) \ge 5-(x+3) \] \[\Large -2x+6 \ge 5-x-3 \] \[\Large -2x+6 \ge -x+2 \] \[\Large -2x \ge -x+2-6 \] \[\Large -2x+x \ge 2-6 \] \[\Large -x \ge 2-6 \] \[\Large -x \ge -4 \] \[\Large x \le \frac{-4}{-1} \] \[\Large x \le 4 \]

OpenStudy (anonymous):

woahhhhhh